OpenStudy (anonymous):

Solve the equation by completing the square. Round to the nearest hundredth if necessary. x2 – 6x = –8 (1 point) 4, –2 –4, 2 4, 2 –4, –2

OpenStudy (anonymous):

\[x^2 – 6x = –8 \implies x^2 – 6x+9=9 –8 \implies (x-3)^2=1 \implies x-3= \pm 1\]\[\implies x=3 \pm 1 \implies x=2~or~x=4\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No sweat. Do math every day.

OpenStudy (anonymous):

I have a few more if you want to help me haha